OpenStudy (anonymous):

A square has a diagonal length of 39 inches. What is the value of any side of the square?

OpenStudy (anonymous):

\[c^2=a^2+b^2\] where c is the diagonal length (39 inches) and a=b since it is a square

OpenStudy (anonymous):

\[c^2=b^2+b^2\]\[39^2=2b^2\]\[1521=2b^2\]\[760.5=b^2\]\[b=27.58\]

OpenStudy (binary3i):

just divide the digonal with square root of 2 and it will give the length of the side

OpenStudy (anonymous):

but what do you mean? (binary3i)

OpenStudy (anonymous):

Zed, first reply is consistent to Pythagoras theorem which is again consistent to the cosine rule ...

OpenStudy (anonymous):

zed's*

OpenStudy (anonymous):

\[\Large \begin{array}{l} {c^2} = {a^2} + {b^2}\\ {a^2} = {b^2}\\ {c^2} = 2{a^2}\\ {a^2} = \frac{{{c^2}}}{2}\\ a = \frac{c}{{\sqrt 2 }} \end{array}\]

OpenStudy (anonymous):

Is what I think binary meant.

OpenStudy (binary3i):

yes

OpenStudy (anonymous):

so the answer is 15?

OpenStudy (anonymous):

Zeds answer was correct 27.58

OpenStudy (anonymous):

my answer choices are 4in, 12in, 15in, and 16in.

OpenStudy (anonymous):

2*4^2 = 32, 2*12^2 = 288, 2*15^2 = 450, 2*16^2 = 512. 39^2 = 1521.

OpenStudy (anonymous):

So either I'm misunderstanding something fundamental, or there's some wrong information here?

OpenStudy (anonymous):

ya wrong info somewhere

OpenStudy (anonymous):

ok..its a square and basically from one corner to another corner its 39 inches. And i need to find the length of any side of the square.

OpenStudy (anonymous):

 Is this right? Where c = 39?